Out of a sample of 80 students from a college 32 were female students determine the 95% confidence interval of the proportion of female Students in the population
1
Expert's answer
2021-07-20T07:16:45-0400

"\hat{p}=\frac{32}{80}=0.4 \\

n=80"

The formula for a CI for a population proportion is

"\hat{p}±z \times \sqrt{ \frac{\hat{p}(1- \hat{p})}{n} }"

z=1.96 for 95 % confidence interval

CI:

"0.4 ±1.96 \times \sqrt{ \frac{0.4(1- 0.4)}{80} } \\

0.4 ± 1.96 \times 0.05477 \\

0.4 ±0.107 \\

(0.293, 0.507)"